# Finite-Time Analysis of Gradient Descent for Shallow Transformers

### v1

@@ 7 CONCLUSION
We performed a finite-time NTK analysis of projected gradient methods for multi-head shallow Transformers. We established nonasymptotic convergence bounds with explicit dependencies on the training period, network width, ambient dimension, and the number of samples. To the best of our knowledge, this is the first work that establishes near-optimality of first-order methods without massive overparameterization for Transformers. Interesting future directions that can build on this work include the study of deep Transformer architectures and the extension of the setting to sequence-to-sequence learning.
@@
In Appendix A, we restate and prove the results in the main text, and in Appendix B, we give further experimental details along with the anonymous code repository to reproduce the experiments.
@@ A.3 Proof of Lemma 2
where $B_{U,m}(\delta^{\prime})$ is defined as in Remark 7.
@@
$\big\|\nabla_{\theta_{i}}^{2}h_{i}\big\|_{F}\;\leq\;\sigma_{2}\underbrace{\big\|\nabla_{\theta_{i}}z_{i}\big\|_{2}^{2}}_{(i)}\;+\;\sigma_{1}\underbrace{\big\|\nabla_{\theta_{i}}^{2}z_{i}\big\|_{F}}_{(ii)}.$ (15)
@@
$D\!\left[\nabla_{U_{i}}z_{i}\right][\Delta]\;=\;D[a_{i}][\Delta]\;=\;M_{i}\,\Delta\,q_{X},\qquad\Rightarrow\qquad\big\|\nabla^{2}_{U_{i},\operatorname{vec}(W_{i})}z_{i}\big\|_{F}\;\leq\;\|M_{i}\|_{F}\,\|q_{X}\|_{2}\;\leq\;1.$
@@
which implies $\big\|\nabla^{2}_{\operatorname{vec}(W_{i})}z_{i}\big\|_{F}\;\leq\;8\,\|U_{i}\|_{2}.$ Combining the $W_{i}$$U_{i}$ and $W_{i}$$W_{i}$ blocks gives
$\big\|\nabla_{\theta_{i}}^{2}z_{i}\big\|_{F}^{2}\;\leq\;2\big\|\nabla^{2}_{U_{i},\operatorname{vec}(W_{i})}z_{i}\big\|_{F}^{2}+\big\|\nabla^{2}_{\operatorname{vec}(W_{i})}z_{i}\big\|_{F}^{2}\;\leq\;2+64\,\|U_{i}\|_{2}^{2},$ (17)
hence $\|\nabla_{\theta_{i}}^{2}z_{i}\|_{F}\leq\sqrt{2+64\|U_{i}\|_{2}^{2}}\leq 8\sqrt{1+\|U_{i}\|_{2}^{2}}$.
@@
$\big\|\nabla_{\theta_{i}}^{2}h(X;\theta_{i})\big\|_{F}\;\leq\;\sigma_{2}\,(1+\|U_{i}\|_{2}^{2})+8\sigma_{1}\,\sqrt{1+\|U_{i}\|_{2}^{2}}.$
@@
$\big\|\nabla_{\theta_{i}}h(X;\theta_{i})\big\|_{2}\leq\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}}\quad\text{and}\quad\big\|\nabla_{\theta_{i}}^{2}h(X;\theta_{i})\big\|_{F}\leq\sigma_{2}\!\big(1+B_{U,m}(\delta^{\prime})^{2}\big)+8\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}}.$
@@ A.5 Proof of Lemma 4
$K_{m}(X,X^{\prime})\;=\;\frac{1}{m}\sum_{i=1}^{m}\!\Biggl[\underbrace{\frac{\partial f}{\partial c_{i}}(X;\varphi^{(0)})\,\frac{\partial f}{\partial c_{i}}(X^{\prime};\varphi^{(0)})}_{\text{$c$–block}}\;+\;\underbrace{\big\langle\nabla_{U_{i}}f(X;\varphi^{(0)}),\,\nabla_{U_{i}}f(X^{\prime};\varphi^{(0)})\big\rangle}_{\text{$U$–block}}\\ \;+\;\underbrace{\big\langle\nabla_{W_{i}}f(X;\varphi^{(0)}),\,\nabla_{W_{i}}f(X^{\prime};\varphi^{(0)})\big\rangle_{F}}_{\text{$W$–block}}\Biggr].$
@@
$\frac{1}{m}\sum_{i=1}^{m}\frac{\partial f}{\partial c_{i}}(X;\varphi^{(0)})\,\frac{\partial f}{\partial c_{i}}(X^{\prime};\varphi^{(0)})$ $=\frac{1}{m}\sum_{i=1}^{m}\sigma\!\big({U_{i}}^{\top}a(X;W_{i})\big)\,\sigma\!\big(U_{i}^{\top}a(X^{\prime};W_{i})\big)$
@@
$\frac{1}{m}\sum_{i=1}^{m}\big\langle\nabla_{U_{i}}f(X;\varphi^{(0)}),\,\nabla_{U_{i}}f(X^{\prime};\varphi^{(0)})\big\rangle$ $=\frac{1}{m}\sum_{i=1}^{m}c_{i}^{2}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\;\langle a_{i},a_{i}^{\prime}\rangle$
@@
Writing $M_{i}=X\,J_{s}(X^{\top}W_{i}q_{X})\,X^{\top}$ and $M_{i}^{\prime}=X^{\prime}\,J_{s}(X^{\prime\top}W_{i}q_{X^{\prime}})\,X^{\prime\top}$, using the Frobenius product identity $\langle xy^{\top},x^{\prime}y^{\prime\top}\rangle_{F}=\langle x,x^{\prime}\rangle\,\langle y,y^{\prime}\rangle$ and that $M$ and $M^{\prime}$ are symmetric,
$\frac{1}{m}\sum_{i=1}^{m}\big\langle\nabla_{W_{i}}f(X),\,\nabla_{W_{i}}f(X^{\prime})\big\rangle_{F}$ $=\frac{1}{m}\sum_{i=1}^{m}\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\big\langle(M_{i}U_{i})q_{X}^{\top},\;(M_{i}^{\prime}U_{i})q_{X^{\prime}}^{\top}\big\rangle_{F}$
@@
$K_{w}(X,X^{\prime})\;:=\;\lim_{m\to\infty}\frac{2}{m}\sum_{i=1}^{m/2}\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\,\langle q_{X},q_{X^{\prime}}\rangle\,U_{i}^{\top}M_{i}M_{i}^{\prime}U_{i}=\langle q_{X},q_{X^{\prime}}\rangle\,\mathbb{E}\!\left[\sigma^{\prime}(U^{\top}a)\sigma^{\prime}(U^{\top}a^{\prime})\,U^{\top}MM^{\prime}U\right].$
@@ A.6 Proof of Lemma 5
$\varepsilon_{\mathrm{app}}\;\leq\;2(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w})\sqrt{\tfrac{2\log(2n/\delta)}{m}}=:B_{\mathrm{app},m}(\delta),$

### v2

@@ 7 CONCLUSION
We performed a finite-time NTK analysis of projected gradient methods for multi-head shallow Transformers. We established nonasymptotic convergence bounds with explicit dependencies on the training period, width, and the number of samples. To the best of our knowledge, this is the first work that establishes near-optimality of first-order methods without massive overparameterization for Transformers. Interesting future directions that can build on this work include the study of deep Transformer architectures and the extension of the setting to sequence-to-sequence learning.
Acknowledgements
We thank the anonymous reviewers for their thoughtful comments and helpful suggestions, which improved the clarity and presentation of the paper. This work was supported in part by the NSF grants: NSF AI Institute (AI-EDGE) 2112471, CNS-NeTS-2106679, the ONR Grant N00014-19-1-2621; and the ARO Grant W911NF-24-1-0103.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] Justification: The setup, the model, and the algorithm are described in Section 3. The function that is assumed to generate the underlying data is described in Section 5.4.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] Justification: The analysis of the algorithms is informally made in Section 4 and formally made in Section 5.5.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes] Justification: The source code is provided at https://github.com/enesarda22/nonasymptotic-transformer.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes] Justification: Every stated theoretical result contains the assumptions if any.
    2. (b)
      Complete proofs of all theoretical results. [Yes] Justification: Complete proofs are provided in the Supplementary Material.
    3. (c)
      Clear explanations of any assumptions. [Yes] Justification: The only assumption we make is on the function generating the underlying data, which is explained in Section 5.4.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes] Justification: The code and instructions are provided in a Github repo in the Supplementary Material.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes] Justification: Experimental details are provided in Section 6 and additional details are provided in the Supplementary Material.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes] Justification: Experimental details are provided in Section 6 and additional details are provided in the Supplementary Material.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [No] Justification: Conducted experiments are not resource-heavy, so we do not describe the computing infrastructure.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Acknowledgements
In Appendix A, we restate and prove the results in the main text, and in Appendix B, we give further experimental details along with the code to reproduce the experiments.
@@ A.3 Proof of Lemma 2
where $B_{U,m}(\delta^{\prime})$ is defined as in (9).
@@
$\big\|\nabla_{\theta_{i}}^{2}h_{i}\big\|_{\mathrm{op}}\;\leq\;\sigma_{2}\underbrace{\big\|\nabla_{\theta_{i}}z_{i}\big\|_{2}^{2}}_{(i)}\;+\;\sigma_{1}\underbrace{\big\|\nabla_{\theta_{i}}^{2}z_{i}\big\|_{\mathrm{op}}}_{(ii)}.$ (16)
@@
$D\!\left[\nabla_{U_{i}}z_{i}\right][\Delta]\;=\;D[a_{i}][\Delta]\;=\;M_{i}\,\Delta\,q_{X},\quad\Rightarrow\quad\big\|\nabla^{2}_{U_{i},\operatorname{vec}(W_{i})}z_{i}\big\|_{\mathrm{op}}\leq\big\|\nabla^{2}_{U_{i},\operatorname{vec}(W_{i})}z_{i}\big\|_{F}\;\leq\;\|M_{i}\|_{F}\,\|q_{X}\|_{2}\;\leq\;1.$
@@
which implies
$\big\|\nabla^{2}_{\operatorname{vec}(W_{i})}z_{i}\big\|_{\mathrm{op}}\;\leq\;8\,\|U_{i}\|_{2}.$
Let $B_{i}:=\nabla^{2}_{U_{i},\operatorname{vec}(W_{i})}z_{i}$ and $C_{i}:=\nabla^{2}_{\operatorname{vec}(W_{i})}z_{i}.$ Then
$\nabla_{\theta_{i}}^{2}z_{i}=\begin{pmatrix}0&B_{i}^{\top}\\ B_{i}&C_{i}\end{pmatrix}$.
For any $(u,v)$ with $\|u\|_{2}^{2}+\|v\|_{2}^{2}=1$,
$\left\|\begin{pmatrix}0&B_{i}^{\top}\\ B_{i}&C_{i}\end{pmatrix}\binom{u}{v}\right\|_{2}^{2}\leq\|B_{i}\|_{\mathrm{op}}^{2}\|v\|_{2}^{2}+\big(\|B_{i}\|_{\mathrm{op}}\|u\|_{2}+\|C_{i}\|_{\mathrm{op}}\|v\|_{2}\big)^{2}\leq 2\|B_{i}\|_{\mathrm{op}}^{2}+\|C_{i}\|_{\mathrm{op}}^{2}.$
Therefore
$\big\|\nabla_{\theta_{i}}^{2}z_{i}\big\|_{\mathrm{op}}\;\leq\;\sqrt{2\big\|\nabla^{2}_{U_{i},\operatorname{vec}(W_{i})}z_{i}\big\|_{\mathrm{op}}^{2}+\big\|\nabla^{2}_{\operatorname{vec}(W_{i})}z_{i}\big\|_{\mathrm{op}}^{2}}\;\leq\;\sqrt{2+64\,\|U_{i}\|_{2}^{2}}\;\leq\;8\sqrt{1+\|U_{i}\|_{2}^{2}}.$ (18)
@@
$\big\|\nabla_{\theta_{i}}^{2}h(X;\theta_{i})\big\|_{\mathrm{op}}\;\leq\;\sigma_{2}\,(1+\|U_{i}\|_{2}^{2})+8\sigma_{1}\,\sqrt{1+\|U_{i}\|_{2}^{2}}.$
@@
$\big\|\nabla_{\theta_{i}}h(X;\theta_{i})\big\|_{2}\leq\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}}\quad\text{and}\quad\big\|\nabla_{\theta_{i}}^{2}h(X;\theta_{i})\big\|_{\mathrm{op}}\leq\sigma_{2}\!\big(1+B_{U,m}(\delta^{\prime})^{2}\big)+8\sigma_{1}\sqrt{1+B_{U,m}(\delta^{\prime})^{2}}.$
@@ A.5 Proof of Lemma 4
$K_{m}(X,X^{\prime})\;=\;\sum_{i=1}^{m}\!\Biggl[\underbrace{\frac{\partial f}{\partial c_{i}}(X;\varphi^{(0)})\,\frac{\partial f}{\partial c_{i}}(X^{\prime};\varphi^{(0)})}_{\text{$c$–block}}\;+\;\underbrace{\big\langle\nabla_{U_{i}}f(X;\varphi^{(0)}),\,\nabla_{U_{i}}f(X^{\prime};\varphi^{(0)})\big\rangle}_{\text{$U$–block}}\\ \;+\;\underbrace{\big\langle\nabla_{W_{i}}f(X;\varphi^{(0)}),\,\nabla_{W_{i}}f(X^{\prime};\varphi^{(0)})\big\rangle_{F}}_{\text{$W$–block}}\Biggr].$
@@
$\sum_{i=1}^{m}\frac{\partial f}{\partial c_{i}}(X;\varphi^{(0)})\,\frac{\partial f}{\partial c_{i}}(X^{\prime};\varphi^{(0)})$ $=\frac{1}{m}\sum_{i=1}^{m}\sigma\!\big({U_{i}}^{\top}a(X;W_{i})\big)\,\sigma\!\big(U_{i}^{\top}a(X^{\prime};W_{i})\big)$
@@
$\sum_{i=1}^{m}\big\langle\nabla_{U_{i}}f(X;\varphi^{(0)}),\,\nabla_{U_{i}}f(X^{\prime};\varphi^{(0)})\big\rangle$ $=\frac{1}{m}\sum_{i=1}^{m}c_{i}^{2}\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\,\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\;\langle a_{i},a_{i}^{\prime}\rangle$
@@
Writing $M_{i}=X\,J_{s}(X^{\top}W_{i}q_{X})\,X^{\top}$ and $M_{i}^{\prime}=X^{\prime}\,J_{s}(X^{\prime\top}W_{i}q_{X^{\prime}})\,X^{\prime\top}$, using the Frobenius product identity $\langle xy^{\top},x^{\prime}y^{\prime\top}\rangle_{F}=\langle x,x^{\prime}\rangle\,\langle y,y^{\prime}\rangle$ and that $M_{i}$ and $M_{i}^{\prime}$ are symmetric,
$\sum_{i=1}^{m}\big\langle\nabla_{W_{i}}f(X),\,\nabla_{W_{i}}f(X^{\prime})\big\rangle_{F}$ $=\frac{1}{m}\sum_{i=1}^{m}\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\big\langle(M_{i}U_{i})q_{X}^{\top},\;(M_{i}^{\prime}U_{i})q_{X^{\prime}}^{\top}\big\rangle_{F}$
@@
$K_{w}(X,X^{\prime})$ $:=\lim_{m\to\infty}\frac{2}{m}\sum_{i=1}^{m/2}\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}\big)\sigma^{\prime}\!\big(U_{i}^{\top}a_{i}^{\prime}\big)\,\langle q_{X},q_{X^{\prime}}\rangle\,U_{i}^{\top}M_{i}M_{i}^{\prime}U_{i}$
$=\langle q_{X},q_{X^{\prime}}\rangle\,\mathbb{E}\!\left[\sigma^{\prime}(U^{\top}a)\sigma^{\prime}(U^{\top}a^{\prime})\,U^{\top}MM^{\prime}U\right].$
@@ A.6 Proof of Lemma 5
Let $\tilde{f}(\cdot;v)\in\mathcal{F}_{\bar{\nu}}$ and $\tilde{\varphi}$ be as in (11). Define
$\varepsilon_{\mathrm{app}}:=\sup_{j\in[n]}\ \big|f_{\mathrm{lin}}(X^{(j)};\tilde{\varphi})-\tilde{f}(X^{(j)};v)\big|.$
Then for any $\delta\in(0,1)$, we have
$\varepsilon_{\mathrm{app}}\;\leq\;4(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w})\sqrt{\tfrac{\log(2n/\delta)}{m}}=:B_{\mathrm{app},m}(\delta),$